OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation x + 2y = 4. A) x-intercept: (0, 2), y-intercept: (4, 0) B) x-intercept: (2, 0), y-intercept: (4, 0) C) x-intercept: (0, 4), y-intercept: (2, 0) D) x-intercept: (4, 0), y-intercept: (0, 2)

OpenStudy (anonymous):

To find x-intercept, let y = 0, and solve for x. To find y-intercept, let x = 0, and solve for y.

OpenStudy (anonymous):

OK so the answered would be wrote like this [9(0)-2(0)=18 is that right

OpenStudy (anonymous):

Or another option: you can solve for y and graph it to see the intercepts. y=-1/2x+2 so D would be the answer

OpenStudy (anonymous):

To solve it how @Easyaspi314 said, you'd have x + 2(0) = 4 x=4 so there's x intercept. 0 + 2y = 4 y=2 y intercept

OpenStudy (anonymous):

Yes, so (4,0) is the x-intercept.

OpenStudy (anonymous):

(0,2) is the y-intercept.

OpenStudy (anonymous):

Choice (D)